# Tackling the Accuracy-Interpretability Trade-off in a Hierarchy of Machine Learning Models for the Prediction of Extreme Heatwaves

### v1

@@ 4.3 A remark on regression and classification
In this work, the different neural network architectures were trained to minimize the CRPS loss, but virtually identical skills can be obtained minimizing the NLL loss. Similarly, the Gaussian approximation minimizes a regression metric (see Equation (6)). On the other hand, training on BCE yields a worse performance (even for the BCES metric itself). Indeed, training on a classification task separates the training data in two classes (heatwave and non-heatwave), neglecting the information about the heatwave amplitude. For the particular problem of heatwaves over France, the choice of the threshold to distinguish the two classes is arbitrary, and there is no obvious regime shift between mild and very extreme heatwaves (Noyelle et al., 2024; Mascolo et al., 2024).
@@ 5.1 GA and IINN
We now consider the projection patterns to discuss the physical interpretation of the predictions (Fig. 1, top panels). These patterns are very robust across the folds and look very similar for GA and IINN, the main difference being a greater relative importance of soil moisture in the former. The most prominent features of the projection patterns are the negative weights for soil moisture, and the large positive coefficients for geopotential over Western Europe. This is in agreement with the understanding that persistent anticyclonic anomalies and dry soil are important predictors of high temperatures in the mid-latitudes (Perkins, 2015; Ardilouze et al., 2017; Weirich-Benet et al., 2023; Barriopedro et al., 2023; Miloshevich et al., 2023a). Note that both models attribute larger weight to soil moisture in Northern France than in Southern France, presumably due to larger variability there making it a slightly better predictor. Another notable property of the projection map for geopotential is the presence of a coherent pattern over the North Atlantic, with positive anomalies over the storm tracks region and negative anomalies over the ocean contributing to the prediction of high temperatures, suggesting anomalies in the path or strength of the Jet Stream. However, the projection map does not allow us to distinguish between different mechanisms pointed out previously, such as stationary Rossby waves (Screen and Simmonds, 2014; Petoukhov et al., 2013; Di Capua et al., 2021; Ragone and Bouchet, 2021; Miloshevich et al., 2023c) or transient eddies (Lehmann and Coumou, 2015).
@@ 5.2 CNN
The previous methods were simple enough that it is possible to relate the trained weights to the final prediction in a rather intuitive way, which allowed us to explain the prediction at once for any input (global interpretability). This is not the case for the CNN, which behaves like a black box and thus is not methodologically interpretable. However, we can still gain physical interpretability by using post-hoc explainable AI techniques, and most will give a different explanation for each input (local interpretability). These methods are quite sensitive to the weights of the neural network, and so when we applied them to the 5 CNN models, we got slightly different results. However, the qualitative conclusions are the same, so, for simplicity, we show the results only for one of the 5 folds.
@@ 5.2.1 Global interpretability: Optimal Input
A possibility to investigate the prediction of the CNN is computing the input (i.e. a map of geopotential height and soil moisture anomalies) that, once fed to the CNN, yields the highest predicted heatwave amplitude. This method, referred to as Backward Optimization (Olah et al., 2017), has been showcased for various problems in weather and climate, such as tornado and hail prediction (McGovern et al., 2019), identification of ENSO phases and seasonal prediction of continental temperatures in North America (Toms et al., 2020). A naïve application of the method consists in initializing our input with a given snapshot from the test data and then performing gradient descent of the loss , not on the weights of the CNN but on the coordinates of . However, doing so yields very noisy maps and absurdly high predicted heatwave amplitudes (see Fig. S6 of the Supplementary Material). Hence, we add regularization terms to constrain the L2 norm and roughness of the optimized input :
@@
where and were empirically chosen as the average L2 norm and roughness of the test data (see Fig. S7 of the Supplementary Material). Similarly, we empirically chose and , which ensured that L2 norm and roughness of the optimized input were close to and , respectively. We then run the optimization initializing with all the snapshots in the test dataset, which gives us slightly different optimized inputs due to the non-convex nature of the optimization. The mean and standard deviation are shown in Fig. 2. We note that in spite of the constraints in the loss, the Z500 optimal input maps do not resemble physical geopotential anomalies, as most of their mass is concentrated over the Atlantic sector. Indeed, the optimization procedure yields heatwave amplitudes which are much larger than any value in the dataset, with an average (see Fig. S8 of the Supplementary Material for the full histogram).
@@
Nevertheless, the average optimal input looks very similar to the projection pattern of the Gaussian approximation (top left panel of Fig. 1), with the notable difference that it attributes a smaller weight to the positive geopotential anomaly over Western Europe and that this anomaly extends further North over Scandinavia. The latter property is reminiscent of the findings of Della-Marta et al. (2007) who report statistical connections between high pressure in this area and heatwaves over Western Europe in historical data. It also hints at a connection to the positive NAO and Atlantic low regimes identified by Cassou et al. (2005). Moreover, in the bottom panel of Fig. 2, we see that most regions of the globe have a very low variance (dark shades), indicating the robustness of the average optimal input pattern. A few regions in the geopotential field, namely the Norwegian Sea and, to a lesser extent, the Canadian and Siberian Arctic, show an amplification of the unitary variance initially present in the seeds (red shades). This may suggest that in these regions the information in the geopotential height field is used non-linearly by the CNN.
@@
Since the GA gives already very good predictions, will be a small non-linear perturbation, and indeed, as can be seen by Fig. S8 of the Supplementary Material, it contributes only for a sixth of the total predicted heatwave amplitude. Now, when we follow the gradient to optimize ,
Importantly, the first term of the right-hand side does not depend on , while the second one will act differently during the optimization as evolves and as different initialization seeds are used. It is thus reasonable that once we take the average over all the test set, this term gets close to 0. However, it will be responsible for the pattern observed in the variance of the optimal input, and thus we can conclude that the sources of non-linear information are mainly concentrated in the Arctic region.
We can make the previous statement more precise by putting our focus directly on , and we can achieve this by adding to the loss another regularization term , which forces the optimization to move in an orthogonal direction with respect to . Doing so yields the results in Fig. 3, with an average .
Concerning the variance, we have in general higher values with respect to Fig. 2, and the focus is more shifted towards the Canadian Arctic. However, the pattern is overall quite similar to the one of Fig. 2. On the other hand, this time the average optimal input pattern highlights new interesting features, the first being a North-East to South-West gradient in soil moisture. A possible explanation is that North-Eastern France is generally wetter, so when that region gets dry, the heatwave will be non-linearly more intense. The patterns observed in the geopotential height field are harder to interpret, with the most prominent feature being a strong anticyclone over the Norwegian Sea, which further reinforces the findings of Della-Marta et al. (2007); Cassou et al. (2005) discussed before. The remaining features of the geopotential height field are not robust across the 5 different models, which underlines the methodological limitations to interpretability beyond linear regression. In the next section we attempt to overcome these limitations with a different method.
Note that we tried performing K-means clustering on the optimal input maps (both with and without the orthogonality constraint), similarly to what is done in Toms et al. (2021). Although the distribution of many individual pixels is markedly bimodal (not shown), the maps overall do not appear to be organized in distinct clusters.
@@ 5.2.2 Local interpretability: Expected Gradients
Consider one input , which is the concatenation of the Z500 and SM fields in our case. Integrated Gradients computes pixel attribution at pixel as
@@
where represents one of the outputs of the neural network for which we seek insight; in our case it could be , , or directly the heatwave probability. is a baseline input, such as the mean input on the training set for instance, which is here zero since our data is normalized. Importantly, the sum of over all pixels should yield the predicted value , according to the completeness axiom (see Sundararajan et al. (2017) for further details). Thus, , which can be either positive or negative, quantifies the contribution of pixel to the predicted output . We used an extension of this method called Expected Gradients (Erion et al., 2021), which reformulates the integral in Eq. 10 as an expectation and combines that expectation with sampling reference values from a background dataset (taken as the entire training set here). We used the implementation provided in the SHAP package (Lundberg and Lee, 2017). Below, we focus on the geopotential height anomaly input for predicting . Results for soil moisture and the prediction of are shown in Figs. S4 and S5 of the Supplementary Material.
@@
To illustrate the feature importance derived from the CNN predictions on relevant cases, we randomly selected examples of heatwaves (with ) from the test dataset. The corresponding Z500 initial conditions are shown in Fig. 4 (first row), together with the resulting Expected Gradients feature importance (EGFI) maps for the CNN prediction (second row). In practice, EGFI maps are smoothed by a Gaussian filter for visual purposes (see Fig. S3 of the Supplementary Material).
@@
All the examples considered exhibit a prominent positive feature over France, which as expected corresponds to the contribution of anticyclonic anomalies to the predicted probability of high temperature over France. In general this feature is the dominant one, but it is not the case for the first example. The feature importance maps also reveal the role of some regions in the North Atlantic and over North America for the CNN prediction. Positive geopotential anomalies in the storm tracks region are consistently associated with positive contribution to predicted temperatures (in particular in examples 1 and 2), as well as negative anomalies over the Labrador sea or Icelandic lows (see in particular example 1). Another region of importance is the Eastern Atlantic Ocean: anticyclonic anomalies off the coast of Portugal contribute negatively to predicted temperatures (examples 1, 2 and 4). This behavior could be related to the role pointed out recently of deep depressions in this region for extreme temperature over Western Europe (Faranda et al., 2023; D’Andrea et al., 2024).
@@
As we are interested in understanding what the CNN has learned beyond the GA, we compared the EGFI maps of the CNN with the EGFI maps of the GA. As the GA model predicts linearly, its EGFI maps are exactly equal to the input multiplied by the GA projection pattern; we show these maps in the third row of Fig. 4. The comparison reveals that the patterns of feature importance are strikingly similar, with the regions of importance largely overlapping. This suggests that the features described above correspond to some regions of the large-scale patterns identified in the projection maps in section 5.1; although no single individual Z500 anomaly snapshot exhibits such a large-scale coherent structure, they all project to some extent on this pattern. Looking at the difference between the EGFI maps of the GA and those of the CNN (Fig. 4, fourth row), we observe that the additional contribution leading to the prediction of the CNN sometimes systematically reinforces the weight in regions already identified by the linear regression performed in the GA (anticyclonic anomaly over France, and to a smaller extent, storm tracks region), sometimes on the contrary, systematically tones down the signal seen in the GA (e.g. Icelandic low), and sometimes has a different sign for different inputs (e.g. off the coast of Portugal). In most cases, this correction also makes the CNN patterns more localized, indicating that, due to its nonlinearity and convolutional nature, the CNN model might be capturing more complex interactions over multiple scales that the GA model, due to its linear nature, cannot represent.
@@
Extracting more detailed information from these plots is challenging because this local input-dependent approach only allows for a qualitative analysis. A global feature importance map could be generated by averaging the absolute values of the local maps over the whole dataset. However, like the optimal input method presented in the previous section, such approaches are limited, as they only highlight the areas of the input deemed important by the CNN model. They do not elucidate how this information is used to make predictions (Rudin, 2019). A particularly relevant question is which scales are significant and how different scales interact. An ad-hoc attempt to address this issue for any black-box model (including CNNs) is discussed in Kasmi et al. (2023). In the next section, we partially address this question with the ScatNet model.
@@ 5.3 ScatNet
At the end of the scattering transform, we obtain a feature map with the shape , where and are the number of latitude and longitude points in the input image, is the number of scales, and is the number of orientations. The first feature map corresponds to the zeroth-order scattering transform, which is just the low-pass-filtered input. The subsequent feature maps correspond to the first-order scattering transform. These maps are generated by convolving the input field with wavelet filters at each scale and orientation (since and ), followed by applying the modulus operator and pooling (Bruna and Mallat, 2013). Similarly to the GA and INN models, we can project the learned weights onto spatial maps for each channel of the feature maps. For the weights to be proper (global) feature importance maps, we normalize them to account for the different ranges of values of each feature map after the scattering transform. Namely, the feature importance for each feature with associated weight in the final linear layer of the network is given by
@@
A natural question is whether we can achieve a predictive performance similar to the GA model with a lower resolution Z500 field. It amounts to know which is the effective scale for a linear prediction. To answer this question, we define ScatNetcoarse, a variant of ScatNet where only the zeroth order features (the coarse Z500 field) are computed, rather than proceeding to the first order in the scattering transform. In Table 4, we compare its predictive performance with the GA and CNN models.
@@
To further validate this statement, Table 5 compares the relative global feature importance for different scales of the Z500 coarse field, and the pixels of SM. These values are derived by calculating the feature-wise mean of the absolute feature importance values across the entire test dataset and then, independently for each scale, summing over geographical locations and different orientations.
@@
Nevertheless, Table 5 indicates that additional information at the sub-synoptic scale ( and below) has an important role for prediction.
Finally, we present in Fig. 6 the feature importance patterns derived from all the first-order scattering features at scale . For scale and , the feature importance maps are very similar, but with less feature importance (see Fig. S10 of the Supplementary Material).

### v2

@@ 4.3 A remark on regression and classification
In this work, the different neural network architectures were trained to minimize the CRPS loss, but virtually identical skills can be obtained minimizing the NLL loss. Similarly, the Gaussian approximation minimizes a regression metric (see Eq. (6)). On the other hand, training on BCE yields a worse performance (even for the BCES metric itself). Indeed, training on a classification task separates the training data in two classes (heatwave and non-heatwave), neglecting the information about the heatwave amplitude. For the particular problem of heatwaves over France, the choice of the threshold to distinguish the two classes is arbitrary, and there is no obvious regime shift between mild and very extreme heatwaves (Noyelle et al., 2024; Mascolo et al., 2024).
@@ 5.1 GA and IINN
We now consider the projection patterns to discuss the physical interpretation of the predictions (Fig. 1, top panels). These patterns are very robust across the folds and look very similar for GA and IINN, the main difference being a greater relative importance of soil moisture in the former. The most prominent features of the projection patterns are the negative weights for soil moisture, and the large positive coefficients for geopotential over Western Europe. This is in agreement with the understanding that persistent anticyclonic anomalies and dry soil are important predictors of high temperatures in the mid-latitudes (Perkins, 2015; Ardilouze et al., 2017; Weirich-Benet et al., 2023; Barriopedro et al., 2023; Miloshevich et al., 2023a). Note that both models attribute larger weight to soil moisture in Northern France than in Southern France, highlighting greater predictability from the former. Another notable property of the projection map for geopotential is the presence of a coherent pattern over the North Atlantic, with positive anomalies over the storm tracks region and negative anomalies over the ocean contributing to the prediction of high temperatures, suggesting anomalies in the path or strength of the Jet Stream. However, the projection map does not allow us to distinguish between different mechanisms pointed out previously, such as stationary Rossby waves (Screen and Simmonds, 2014; Petoukhov et al., 2013; Di Capua et al., 2021; Ragone and Bouchet, 2021; Miloshevich et al., 2023c) or transient eddies (Lehmann and Coumou, 2015).
@@ 5.2 CNN
The previous methods were simple enough that it is possible to relate the trained weights to the final prediction in a rather intuitive way, which allowed us to explain the prediction at once for any input (global interpretability). This is not the case for the CNN, which behaves like a black box and thus is not methodologically interpretable. However, we can still gain physical interpretability by using post-hoc explainable AI techniques, and most will give a different explanation for each input (local interpretability). These methods are slightly sensitive to the weights of the neural network, and so when we applied them to the 5 CNN models, we obtained subtly different results. However, the qualitative conclusions are the same, so, for simplicity, we show the results only for one of the 5 folds.
@@ 5.2.1 Global interpretability: Optimal Input
A possibility to investigate the prediction of the CNN is computing the input (i.e. a map of geopotential height and soil moisture anomalies) that, once fed to the CNN, yields the highest predicted heatwave amplitude. This method, referred to as Backward Optimization (Olah et al., 2017), has been showcased for various problems in weather and climate, such as tornado and hail prediction (McGovern et al., 2019) or identification of ENSO phases and seasonal prediction of continental temperatures in North America (Toms et al., 2020). A naïve application of the method consists in initializing our input with a given snapshot from the test data and then performing gradient descent of the loss , not on the weights of the CNN but on the coordinates of . However, doing so yields very noisy maps and absurdly high predicted heatwave amplitudes (see Fig. S7 of the Supplementary Materials). Hence, we add regularization terms to constrain the L2 norm and roughness of the optimized input :
@@
where and were empirically chosen as the average L2 norm and roughness of the test data (see Fig. S8 of the Supplementary Materials). We searched logarithmically spaced values of the two regularization coefficients, settling on and , which were the smallest values ensuring that L2 norm and roughness of the optimized input were close to and , respectively. We then run the optimization initializing with all the snapshots in the test dataset, which gives us slightly different optimized inputs due to the non-convex nature of the optimization. In the top row of Fig. 2, we show the average optimized input, and we note that in spite of the constraints in the loss, the Z500 optimal input map does not resemble physical geopotential anomalies, as most of the mass is concentrated over the Atlantic sector. Indeed, the optimization procedure yields heatwave amplitudes which are much larger than any value in the dataset, with an average (see Fig. S9 of the Supplementary Materials for the full histogram).
@@
Nevertheless, the average optimal input looks very similar to the projection pattern of the Gaussian approximation (top left panel of Fig. 1), with the notable difference that it attributes a smaller weight to the positive geopotential anomaly over Western Europe and that this anomaly extends further North over Scandinavia. The latter property is reminiscent of the findings of Della-Marta et al. (2007) who report statistical connections between high pressure in this area and heatwaves over Western Europe in historical data. It also hints at a connection to the positive NAO and Atlantic low regimes identified by Cassou et al. (2005).
In the bottom panel of Fig. 2, we plot the signal-to-noise-ratio (STNR) defined as the pixelwise ratio between mean and standard deviation of the optimal inputs (see Fig. S11 of the Supplementary Materials). The very high values of STNR indicate the robustness of the average optimal input pattern. In particular, there is a very strong convergence (i.e. high STNR) of the optimized inputs over Europe and Northern Africa, hinting at the importance that the main anticyclone doesn’t extend too far South but rather stays concentrated over Central Europe. In the teleconnection patterns over the Atlantic, North America and Siberia, there is also a quite strong convergence around thin zonal lines, but this may be an artifact due to the architecture of the CNN. On the other hand, the Arctic region displays particularly small values of the STNR, due to an amplification of the original unitary variance initially present in the seeds (see Fig. S11 of the Supplementary Materials). This may suggest that in this region the information in the geopotential height field is used non-linearly by the CNN.
Concerning soil moisture, the STNR is greatest in the Northeastern corner of France, despite the maximum in the mean being rather in the center, but this is not a robust pattern across the different folds.
@@
Since the GA gives already very good predictions, will be a small non-linear perturbation, and indeed, as can be seen by Fig. S9 of the Supplementary Materials, it contributes only a sixth of the total predicted heatwave amplitude. Now, when we follow the gradient to optimize ,
Importantly, the first term of the right-hand side does not depend on , while the second one will act differently during the optimization as evolves and as different initialization seeds are used. It is thus reasonable that once we take the average over all the test set, this second term gets close to 0. However, it will be responsible for the pattern observed in the variance of the optimal input, and thus it draws our attention to the Arctic.
We can investigate the sources of non-linear information more precisely by putting our focus directly on , and we can achieve this by adding to the loss another regularization term , which strongly nudges the optimization in an orthogonal direction with respect to . Similarly to what we did for and , we chose , which yields the results in Fig. 3, with an average .
This time, the average optimal input pattern highlights new interesting features, the first being a North-East to South-West gradient in soil moisture. A possible explanation is that North-Eastern France is generally wetter, so when that region gets dry, the heatwave will be non-linearly more intense (see Eq. (9)). The patterns observed in the geopotential height field are harder to interpret, with the most prominent feature being a strong anticyclone over the Norwegian Sea, which further reinforces the findings of Della-Marta et al. (2007); Cassou et al. (2005) discussed before. Nevertheless, the signal-to-noise-ratio is quite low everywhere, showing that these second order patterns are not very robust.
Another option to investigate non-linear patterns is to perform K-means clustering on the optimal input maps (both with and without the orthogonality constraint), similarly to what is done in Toms et al. (2021). Although the distribution of many individual pixels is markedly bimodal (not shown), the maps overall do not appear to be organized in distinct clusters.
In the end, obtaining global interpretability beyond the linear model proved extremely challenging, so, in the next section, we focus instead on local interpretability.
@@ 5.2.2 Local interpretability: Expected Gradients
Consider one input , which in our case is the stack of Z500 and SM fields. Integrated Gradients computes pixel attribution at pixel as
@@
where represents one of the outputs of the neural network for which we seek insight; in our case it could be , , or directly the heatwave probability. is a baseline input, such as the mean input on the training set for instance, which is here zero since our data is normalized. Importantly, the sum of over all pixels should yield the predicted value , according to the completeness axiom (see Sundararajan et al. (2017) for further details). Thus, , which can be either positive or negative, quantifies the contribution of pixel to the predicted output . We used an extension of this method called Expected Gradients (Erion et al., 2021), which reformulates the integral in Eq. 10 as an expectation and combines that expectation with sampling reference values from a background dataset (taken as the entire training set here). We used the implementation provided by the method GradientExplainer in the SHAP package (Lundberg and Lee, 2017). Below, we focus on the geopotential height anomaly input for predicting . Results for soil moisture and the prediction of are shown in Figs. S5 and S6 of the Supplementary Materials.
@@
To illustrate the feature importance derived from the CNN predictions on relevant cases, we randomly selected four examples from the 100 most extreme heatwaves in the test set. The corresponding Z500 initial conditions are shown in Fig. 4 (first row), together with the resulting Expected Gradients feature importance (EGFI) maps for the CNN prediction (second row). In practice, EGFI maps are smoothed by a Gaussian filter for visual purposes (see Fig. S3 of the Supplementary Materials).
@@
For such extreme events, the heatwave amplitude was already above the 5% threshold for several days, and the main dynamics is that of an anticyclone over Western Europe progressively getting stronger (see for instance Zschenderlein et al., 2019). It is thus natural that most examples (among the ones shown all but the first) exhibit a prominent positive feature over France. The feature importance maps also reveal the role of some regions in the North Atlantic and over North America for the CNN prediction. Positive geopotential anomalies in the storm tracks region are consistently associated with positive contribution to predicted temperatures (in particular in examples 1 and 2), as well as negative anomalies over the Labrador sea or Icelandic lows (see in particular example 1). Another region of importance is the Eastern Atlantic Ocean: anticyclonic anomalies off the coast of Portugal contribute negatively to predicted temperatures (examples 1, 2 and 4). This behavior could be related to the recently highlighted role of deep depressions in this region for extreme temperature over Western Europe (Faranda et al., 2023; D’Andrea et al., 2024).
As we are interested in understanding what the CNN has learned beyond the GA, we compared the EGFI maps of the CNN with the EGFI maps of the GA. As the GA model predicts linearly, its EGFI maps are exactly equal to the input multiplied by the GA projection pattern ; we show these maps in the third row of Fig. 4. The comparison reveals that the patterns of feature importance are strikingly similar, with the regions of importance largely overlapping.
@@
This suggests that the features described above correspond to some regions of the large-scale patterns identified in the projection maps in section 5.1; although no single individual Z500 anomaly snapshot exhibits such a large-scale coherent structure, they all project to some extent on this pattern. Looking at the difference between the EGFI maps of the GA and those of the CNN (Fig. 4, fourth row), we observe that the additional contribution leading to the prediction of the CNN behaves differently in different geographical locations. For instance, over France, and to a lesser extent in the storm tracks region, the CNN systematically reinforces the weight already identified by the GA. In other regions, like the Icelandic low, it systematically tones down the signal seen in the GA. And finally, other locations, like off the coast of Portugal, show a different sign depending on the particular input. In most cases, this correction that the CNN makes on the GA baseline yields more localized CNN patterns. This suggests that, due to its nonlinearity and convolutional nature, the CNN model might be capturing more complex interactions over multiple scales, that the GA model, due to its linear nature, cannot represent.
While extracting detailed information from these plots is challenging due to the qualitative nature of this local input-dependent approach, we have quantified the correlation between the EGFI maps and the GA maps more systematically for all inputs in the test. We obtained an average correlation of (full histogram in Fig. S4 of the Supplementary Materials), which confirms the fact that the EGFI maps of GA and CNN are indeed consistently similar, not only for the very extreme events shown in Fig 4.
@@
Additionally, a global feature importance map can be generated by averaging the absolute values of the local maps across the entire dataset. However, like the optimal input method presented in the previous section, this approach would be limited, as it would only highlight the areas of the input deemed important by the CNN model, not how this information is used to make predictions (Rudin, 2019).
A particularly relevant question is which scales are significant and how different scales interact. An ad-hoc attempt to address this issue for any black-box model (including CNNs) is discussed in Kasmi et al. (2023). In the next section, we partially address this question with the ScatNet model.
@@ 5.3 ScatNet
At the end of the scattering transform, we obtain a feature map with the shape , where and are the number of latitude and longitude points in the input image, is the number of scales, and is the number of orientations. The first feature map corresponds to the zeroth-order scattering transform, which is just the low-pass-filtered input. The subsequent feature maps correspond to the first-order scattering transform. These maps are generated by convolving the input field with wavelet filters at each scale and orientation (since and ), followed by applying the modulus operator and pooling (Bruna and Mallat, 2013). Similarly to the GA and IINN models, we can project the learned weights onto spatial maps for each channel of the feature maps. For the weights to be proper (global) feature importance maps, we normalize them to account for the different ranges of values of each feature map after the scattering transform. Namely, the feature importance for each feature with associated weight in the final linear layer of the network is given by
@@
A natural question is whether we can achieve a predictive performance similar to the GA model with a lower resolution Z500 field. It amounts to know which is the effective scale for a linear prediction. To answer this question, we define ScatNetcoarse, a variant of ScatNet where only the zeroth order features (the coarse Z500 field) are computed, rather than proceeding to the first order in the scattering transform. In Tab. 4, we compare its predictive performance with the GA and CNN models.
@@
To further validate this statement, Tab. 5 compares the relative global feature importance for different scales of the Z500 coarse field, and the pixels of SM. These values are derived by calculating the feature-wise mean of the absolute feature importance values across the entire test dataset and then, independently for each scale, summing over geographical locations and different orientations.
@@
Nevertheless, Tab. 5 indicates that additional information at the sub-synoptic scale ( and below) has an important role for prediction.
Finally, we present in Fig. 6 the feature importance patterns derived from all the first-order scattering features at scale . For scale and , the feature importance maps are very similar, but with less feature importance (see Fig. S13 of the Supplementary Materials).